3-46. Determine the stretch in each of the two springs required to hold the 20-kg crate in the equilibrium position shown. Each spring has an unstretched length of 2 m and a stiffness of k = 300 N/m. for 12 m Prob. 3-406
Answer #1

3.46

FOC= FOC(6i+4j+12k/sqrt(6^2+4^2+12^2))=(3/7)FOCi+(2/7)FOCj+(6/7)FOCk

FOA=-FOAj

FOB=-FOBi

F=-196.2K

SumF=0

Equating I,j,k components

(3/7)FOC-FOB=0

(2/7)FOC-FOA=0

(6/7)FOC-196.2=0

gives,

FOC=228.9N

FOB= 98.1N

FOA=65.4N

By Hooke’s law F=kx

x= F/k

xOB=98.1/300=0.327m

xOA=65.4/300=0.218m